Calculation Exercise

Remember show all calculations and round to two decimals.

Peter Cheng from Cheng Enterprises is conducting an inventory analysis on one of his most popular products. The annual demand for this product is 140 000 units, price per unit is R450, carrying cost is 20% of the unit price, and ordering costs amount to R30 per order. Orders must be placed in round lots of 10. The lead time is 4 (four) days. Assume Peter does business for 48 weeks per year, 5 days per week, and requires a safety stock inventory level of 12 100 units.

Given the above-mentioned information, calculate the following:

2.4     Re-order point.                                                                                                                          (4)
Answer #1

Answer:

  • Total no. of days in a year = no. of weeks * no. days per week = A = 48*5 = 240 days
  • Annual demand = B= 140000
  • Daily demand = C = B/A = 140000 / 240 = 583.33 units
  • Demand during lead time =D = C * lead time = 583.33*4 = 2333.33 units
  • Safety stock = E = 12100
  • Reorder point = D+E = 2333.33+12000= 14333.33 (Answer)